Mr. S. is admitted to the hospital with MRSA abscess of the left foot. His physician...

  1. Mr. S. is admitted to the hospital with MRSA abscess of the left foot. His physician orders Vancomycin IV.   Order: Vancomycin 125 mg in 500 ml D5W to infuse over 2 hours via infusion pump. What is the flow rate in ml/hr?
  1. Mrs. N. suffered a 3rd degree burn to her right hand and was admitted to the hospital for wound debridement and IV antibiotic therapy. Order: Gentamycin 40 mg IV every 8 hours. Supply: Gentamycin 80 mg per 2 ml. How many ml will you give?
  1. Mrs. B. is admitted to the hospital from a Skilled Nursing Facility with failure to thrive and dehydration. Her physician orders Lactated Ringers 1000 ml with multi-vitamin added to infuse over the next 6 hours via infusion pump. What is the flow rate in ml/hr?

  1. Mrs. Y. is admitted to the Critical Care Unit with an acute myocardial infarction. Her physician orders a Lidocaine drip via infusion pump. Order: Lidocaine 2 g in 500 ml D5W IV to infuse at 6 mg/min. What is the infusion rate in ml/min?

  1. Mr. S. is status post appendectomy and is admitted for observation secondary to nausea and vomiting post-operatively. His physician orders NS 1000 ml IV for 12 hours. Drop factor: 20gtt/ml. What is the flow rate in gtt/min?

  1. Mr. A. is admitted to the medical unit with magnesium deficiency. His physician orders a magnesium infusion. Order: Magnesium sulfate 4 g in 500 ml D5W at 500 mg/hr via infusion pump. What is the flow rate in ml/hr?

  1. Mr. M. is admitted to the hospital with a GI bleed and his physician orders a blood transfusion of packed red blood cells (PRBCs). Order: 1 unit PRBCs; 520 ml IV to infuse over 4 hours. Drop factor: 20 gtt/ml. What is the infusion rate in gtt/min?

  1. Mr. H. is admitted to the orthopedic floor after a hip replacement. His physician orders IV fluids to infuse at 150 ml/hr for the next 24 hours. Per this order, how many liters of fluid will infuse?

  1. Mr. O. is admitted to the ICU with cardiac decompensation. His physician orders a Dobutamine drip via infusion pump. Order: Dobutamine 250 mg IV in 500 ml D5W to infuse at 5 mcg/kg/min. Mr. O. weighs 80 kg. What is the flow rate in ml/hr?

  1. Mr. G. is admitted to the hospital with acute pancreatitis and rates his abdominal pain 8/10. His physician orders Hydromorphone PRN for pain management. Order: Hydromorphone 1 mg IV every 4 hours PRN pain. Supply: Hydromorphone 5 mg/ml. How many ml of medication will Mr. G. receive with each dose?
Answer #1

●Given

Order:Vancomycin

Dose:125mg

Volume:500ml

Time:2hours

Flow rate=Total volume (ml)÷Time (hr)

=500÷2

=250

Ans:250ml/hr

●Given

Order:Gentamicin

Dose:40mg

Frequency:8 hourly (three time a day)

Available:80mg/2ml

When 80mg=2ml

The 40mg=??ml

Cross multiply and divide

=40×2÷80

=80÷80

=1

Ans:1ml to be given

●Given

Order:RL with multivitamin

Volume:1000ml

Time:6 hours

Flow rate=Total volume (ml)÷Time (hr)

=1000÷6

=166.6 or 167

Ans:167 ml has to be administered

●Given

Order:Lidocaine

Dose:2gm

Volume:500mlD5W

Drop factor:15 (standard rate)

Infusion rate:6mg/min

If in 1 min 6mg will be transfused

Then

In 60 minutes =??mg

Cross multiply and divide

=60×6÷1

=360÷1

=360mg/hour

Then 2gram(2000mg) will be transfused in

=2000×1÷360

=2000÷360

=5.5 hour(330min =5.5×60)

Flow rate:Total volume (ml)÷Time(min)×drop factor

=500÷330×15

=1.51×15

=22.7 or 23

Ans:23ml/min